Suppose there are two blocks connected bya maeeles sir and put over a pulley. IF the pulley is NOT massless will the tension in the strings be the same? Why not(please prove)

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

As the pulley has also some mass so the tension in both the strings will be different..

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

if the pulley is not mass less then rotational motion of pulley is to be taken in account...

for block 1

             M1g-T1=Ma ...............1

for block 2

             T2-M2g=M2a....................2

for pulley

         (T1-T2)R =I(alfa) =Ia/R

         T1-T2= Ia/R2 ..................3

 here I is the moment of inertial of pulley .....

from these equations we will get different values of tension.......